# Self-Supervised Multisensory Pretraining for Contact-Rich Robot Reinforcement Learning

### v1

@@ Abstract
Effective contact-rich manipulation requires robots to synergistically leverage vision, force, and proprioception. However, Reinforcement Learning agents struggle to learn in such multisensory settings, especially amidst sensory noise and dynamic changes. We propose MultiSensory Dynamic Pretraining (MSDP), a novel framework for learning expressive multisensory representations tailored for task-oriented policy learning. MSDP is based on masked autoencoding and trains a transformer-based encoder by reconstructing multisensory observations from only a subset of sensor embeddings, leading to cross-modal prediction and sensor fusion. For downstream policy learning, we introduce a novel asymmetric architecture, where a cross-attention mechanism allows the critic to extract dynamic, task-specific features from the frozen embeddings, while the actor receives a stable pooled representation to guide its actions. Our method demonstrates accelerated learning and robust performance under diverse perturbations, including sensor noise, and changes in object dynamics. Evaluations in multiple challenging, contact-rich robot manipulation tasks in simulation and the real world showcase the effectiveness of MSDP. Our approach exhibits strong robustness to perturbations and achieves high success rates on the real robot with as few as 6,000 online interactions, offering a simple yet powerful solution for complex multisensory robotic control.
@@ I Introduction
Reinforcement Learning (RL) has shown impressive successes in learning complex tasks ranging from Atari [mnih_playing_2013], locomotion [zhuang_robot_2023], vision-based manipulation [zhang_towards_2015] to multisensory peg insertion [lee_making_2018, sferrazza_power_2023]. However, incorporating multiple sensor modalities, especially in complex contact-rich robotic manipulation tasks, remains a challenge for RL, due to the heterogeneous dynamics of different sensor modalities. Additionally, the importance of each input modality changes during the execution of a manipulation task, e.g., coarse scene understanding from visual input to fine-grained force feedback when in contact. Thus, robotic agents need to learn how to dynamically focus on the most relevant sensory information while adapting to perturbations and dynamic changes in the environment. This capability, referred to as sensor fusion, is a long-studied problem in the field of robotics for various tasks ranging from control [garcia2008sensor, khalil2010dexterous], manipulation [hu2016development, xia2022review], localization, and navigation [alatise2020review], but remains underexplored in RL settings.
To this end, Imitation Learning approaches [li_see_2022, hao_masked_2023, saxena_mrest_2024] have shown promise in utilizing multisensory data for learning skills, but require experts collecting informative data, thus limiting their usability, robustness and generalization. The challenge of data collection becomes particularly pronounced in tasks involving objects with varying or uncertain properties, such as mass or friction. In such cases, substantial and often expensive data acquisition efforts are required to develop a robust policy capable of generalizing effectively across diverse contextual variations [belkhale_dataquality_2024]. Self-directed exploration in RL, on the other hand, facilitates the development of adaptable strategies that can be generalized across diverse object properties and contexts. These advantages position RL as a compelling approach for learning multisensory contact-rich manipulation [lee_making_2018, chen_visuo-tactile_2022, liu_masked_2024]. Recent advances in multimodal learning [bachmann_multimae_2022, geng_multimodal_2022] have shown the advantages of reconstructing masked or noised inputs to learn expressive cross-modal representations for downstream tasks. Such masking-based self-supervision approaches, whether at the sensory or embedding level, have also been shown to improve the network’s robustness [skand_simple_2024, liu_learning_2017, radosavovic_real-world_2022].
@@
To summarize, the key contributions of our work are twofold: (i) we develop an effective pretraining strategy based on masked autoencoding to form a rich multisensory representation, and (ii) we introduce a novel architecture that allows task-specific feature extraction to achieve efficient (real) robot Reinforcement Learning. Our findings pave the way for more adaptive and resilient RL agents to handle multiple input modalities to master complex contact-rich manipulation tasks.
@@ II-A Reinforcement Learning for Contact-Rich Manipulation
Tasks where a robotic manipulator has to interact with its environment, either directly or indirectly via a tool, require a good understanding of the interaction forces that shape the task at hand. Solving contact-rich tasks often relies on an accurate estimate of force and dynamics. When such estimates are available, classical control approaches can be adapted to solve the task [whitney1987historical, whitney1982quasi]. To better handle unseen or difficult contact dynamics in unstructured environments, RL presents itself as an ideal candidate that can learn directly via interactions [nuttin1997learning], without accurate state estimation [levine2016end]. In the context of contact-rich manipulation, various works have employed RL to learn contact-based policies on a variety of tasks [elguea2023review, suomalainen2022survey, liu2021deep].
@@ II-B Multimodal Self-Supervision for Robotics
Lee et al. [lee_making_2018] learns a multimodal representation from vision, force torque, and proprioception via MLP-fusion and multiple self-supervised objectives. The frozen representation leads to a robust representation for RL to solve multiple peg insertion tasks. [chen_visuo-tactile_2022] extended the Vision Transformer [dosovitskiy_image_2020] with a force torque sensor to solve a variety of contact-rich tasks. They additionally incorporate the self-supervised objectives from [lee_making_2018] to shape a representation using SLAC [lee_stochastic_2020]. [mejia_hearing_2024] pretrains an audio-encoder [morgado_audio-visual_2021] to combine vision and audio via a transformer decoder for Imitation Learning. The audio signal from the contact microphone provides rich feedback for various manipulation tasks. To also account for different sensor frequencies [saxena_mrest_2024] developed a multi-resolution policy based on pretrained Vision Language Models to improve inference time using proprioception and force.
@@ II-C Deep Sensor Fusion
A robust representation is essential to integrate and process various sensory inputs to ensure stable and efficient learning in multisensory RL. Previous architectures often focused on straightforward latent fusion approaches by concatenating the various representations for downstream tasks [lee2020making, li2023see, li2021reinforcement, guzey2023dexterity, han2024learning]. Feng et al. [feng2024play] take this a step further by adding a subgoal-aware weighting for learning the stage-wise importance of difference sensors. Alternatively, contrastive learning approaches [dave_multimodal_2024, lygerakis_m2curl_2024] emphasize feature alignment, rather than fusion, to learn a shared latent representation between multiple modalities. However, distilling task-relevant features while maintaining sensor-specific information can be challenging. In contrast, inspired by the success of masked token prediction [bachmann_multimae_2022], recent works have also explored how masked multisensory pretraining can enhance representation learning for contact-rich manipulation [sferrazza_power_2023, liu_masked_2024].
@@ III Preliminaries
We define a multisensorial POMDP as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{O}^{MS},\ \mathcal{A},\ \mathcal{R},\ \mathcal{P},\ \gamma\rangle$ [kurniawati_partially_2021], where $\mathcal{S}$ is the true state of the environment. The agent does not have access to $\mathcal{S}$ and observes it through multiple sensor observations, which we specify as Proprioception $\mathcal{O}^{P}\in\mathbb{R}^{14}$, Force Torque (FT) $\mathcal{O}^{FT}\in\mathbb{R}^{4\times 6}$ and Vision $O^{V}\in\mathbb{R}^{64\times 64\times 3}$. The sensors build up the multisensory Observation-space $\mathcal{O}^{MS}=[\mathcal{O}^{P},\ \mathcal{O}^{FT},\ \mathcal{O}^{V}]$. $\mathcal{A}$ is the action space, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the transition kernel, and $\gamma\in[0,1)$ is the discount factor. Our goal is to learn a policy $\pi:\mathcal{O}^{MS}\to\mathcal{A}$ that maximizes the discounted sum of rewards $\mathbb{E}_{\pi}[\sum_{t}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$.
We aim to extract an expressive, fused representation from all sensors in order to solve contact-rich manipulation tasks with RL. We only provide the current multisensory observation and no history. We use the off-policy algorithm SAC [haarnoja_soft_2019] in simulation and RLPD [ball_efficient_2023] for efficiency in the real world, however we highlight that our approach is compatible with any actor-critic algorithm.
@@ IV Multisensory Dynamic Pretraining
We present MultiSensory Dynamic Pretraining (MSDP), a novel framework for representation learning that builds upon Masked Autoencoders [he_masked_2021], tailored to enhance contact-rich robot reinforcement learning tasks that require perception through multiple sensor modalities. MSDP introduces a modular architecture to seamlessly handle diverse input sensors, and a multisensory masking scheme to promote rich cross-modal representations, i.e., to retain knowledge about the task and the environment dynamics even in the absence or perturbation of one or more modalities.
Furthermore, we shed light on the different ways to map pretrained transformer embeddings to input states for downstream RL tasks, referred to as latent bridging, a problem often overlooked in practice by the community. In this context, we propose employing a simple cross-attention layer to obtain expressive task-specific features from frozen multisensory embeddings for the critic representation, which is also referred to as attention pooling [chen_context_2023]. The actor on the other hand, receives a special pooling of the sensor embeddings to ensure a stable representation. Both mechanisms offer a fixed low-dimensional observation for the downstream RL agent.
@@ IV-A MSDP Architecture
Given the inputs from multiple sensor modalities, such as vision, force torque readings and proprioception, we encode each modality using a separate network referred to as sensor encoders. Particularly for the vision input, we use a CNN-stem similar to [seo_masked_2023], which allows us to mask at embedding- rather than pixel-level, as masked object positions cannot be retrieved in pixel-space. The CNN-stem introduces redundancy as vision embeddings have overlapping receptive fields and, most importantly, stabilizes training [xiao_early_2021]. Furthermore, compared to a patchify-stem, it reduces the burden of the multisensory encoder to extract vision features, while fusing with other modalities. We use a linear projection to encode force torque readings and proprioception given their low dimensionalities.
Subsequently, we incorporate position encoding for the embedded sensor features to contextualize the position and modality of each embedding. Rather than using fixed positional/modality encoding [vaswani_attention_2017] we use learnable embeddings to encode both together. Once we obtain embedding representations from all sensors, we randomly mask out a subset of sensor tokens and feed the remaining to our transformer encoder, resulting in multisensory latent embeddings. The attention mechanism of the encoder leads to dynamic multisensory fusion promoted by the representation objective described in Section IV-B. A learnable mask embedding is added to the randomly masked out embedding positions, next to the multisensory embeddings from the encoder, to generate the original number of embeddings. We reapply our positional/modality encoding before feeding all embeddings to the decoder. During decoding, embeddings exchange information, especially the mask embedding, before being fed to separate decoder heads to fulfill the representation learning objective. We use a shared linear projection for all vision embeddings to reconstruct their corresponding patch.
@@ IV-B Representation Learning
Our representation learning objective is based on multimodal masked autoencoding [he_masked_2021, bachmann_multimae_2022, geng_multimodal_2022]. The objective is to either reconstruct the current $\mathbf{O}^{MS}_{t}$ or next observation $\mathbf{O}^{MS}_{t+1}$ from a random subset of sensor embeddings. Vision has a high number of embeddings and needs to extract information about other, potentially masked, sensors e.g. identifiying contact to estimate force. Non-vision sensors on the other hand, are beneficial to reconstruct the vision observation as e.g. proprioception defines the robot position. Vision as a global sensing modality never gets fully masked out, whereas low-dimensional sensors are not available when masked. This representation objective results in cross-sensor prediction thus leading to fusion of all modalities. Furthermore, the decoder is conditioned on action in the prediction objective and needs to extract dynamic action-related features important for downstream RL. We denote our method MSDP-P (Prediction) wherein we reconstruct the next observation $\mathbf{O}^{MS}_{t+1}-\Phi(\mathbf{O}^{MS}_{t},\mathbf{A}_{t})||^{2}$ and MSDP-R (Reconstruction) where we reconstruct the current observation $\mathbf{O}^{MS}_{t}-\Phi(\mathbf{O}^{MS}_{t})||^{2}$. $\Phi(.)$ denotes the network prediction conditioned on the current observation $\mathbf{O}^{MS}_{t}$ and action $\mathbf{A}_{t}$. $||\cdot||^{2}$ denotes the Mean Squared Error. Sensor embeddings are only masked during pretraining.
@@ IV-C Policy Learning
From the pretrained encoder, we receive expressive multisensory embeddings given the available sensors. Extracting a compact representation from those embeddings for downstream task-solving is a crucial and often under-discussed design choice. We name the mapping between embeddings and a compact representation latent bridging, which has a considerable impact on performance. To address this, we depart from works that naively extract the "CLS"-token from the high-dimensional embeddings, analogous to the Vision Transformer [radosavovic_real-world_2022, dosovitskiy_image_2020, xiao_masked_2022, majumdar_where_2024] and, instead, propose an asymmetric latent bridging strategy between actor and critic to obtain a compact representation. The critic uses a single cross-attention layer with a learnable query and the multisensory embeddings from the MSDP encoder as keys and values (see Figure 2). It offers dynamic task-specific feature extraction (e.g. object positions, robot state, contact) over the task-solving process. Fine-grained understanding of the environment leads to faster convergence compared to a global representation. The policy, on the other hand, does not profit from a cross-attention layer as it may destabilize training. Instead, we mean pool all embeddings originating from the vision sensor before pooling all embeddings, to account for the uneven number of sensor embeddings. Pooling sensor tokens results in a stable and parameter-free latent bridging similiar to [sferrazza_power_2023, seo_masked_2023]. This assymmetric representation between actor and critic, follows [garcin_studying_2025], where the actor benefits from a stable representation over task stages and the critic from a dynamic-specific representation. The cross-attention layer is trained by the critic, resulting in task-specific feature extraction. We train with the off-policy algorithm SAC in simulation and RLPD in the real world, where we incorporate offline data for pretraining in our replay buffer. Our approach is working with any actor-critic RL algorithm.

### v2

@@ Abstract
Effective contact-rich manipulation requires robots to synergistically leverage vision, force, and proprioception. However, Reinforcement Learning agents struggle to learn in such multisensory settings, especially amidst sensory noise and dynamic changes. We propose MultiSensory Dynamic Pretraining (MSDP), a novel framework for learning expressive multisensory representations tailored for task-oriented policy learning. MSDP is based on masked autoencoding and trains a transformer-based encoder by reconstructing multisensory observations from only a subset of sensor embeddings, leading to cross-modal prediction and sensor fusion. For downstream policy learning, we introduce a novel asymmetric architecture, where a cross-attention mechanism allows the critic to extract dynamic, task-specific features from the frozen embeddings, while the actor receives a stable pooled representation to guide its actions. Our method demonstrates accelerated learning and robust performance under diverse perturbations, including sensor noise, and changes in object dynamics. Evaluations in multiple challenging, contact-rich robot manipulation tasks in simulation and the real world showcase the effectiveness of MSDP. Our approach exhibits strong robustness to perturbations and achieves high success rates on the real robot with as few as 6,000 online interactions, offering a simple yet powerful solution for complex multisensory robotic control. Website: https://msdp-pearl.github.io/
@@ I Introduction
Reinforcement Learning (RL) has shown impressive successes in learning complex tasks ranging from Atari [38], locomotion [54], vision-based manipulation [53] to multisensory peg insertion [26, 44]. However, incorporating multiple sensor modalities, especially in complex contact-rich robotic manipulation tasks, remains a challenge for RL, due to the heterogeneous dynamics of different sensor modalities. Additionally, the importance of each input modality changes during the execution of a manipulation task, e.g., coarse scene understanding from visual input to fine-grained force feedback when in contact. Thus, robotic agents need to learn how to dynamically focus on the most relevant sensory information while adapting to perturbations and dynamic changes in the environment. This field of sensor fusion is a long-studied problem in the field of robotics for various tasks ranging from control [13, 22], manipulation [21, 50], localization, and navigation [1], but remains underexplored in RL settings.
To this end, Imitation Learning approaches [30, 19, 42] have shown promise in utilizing multisensory data for learning skills, but require experts collecting informative data, thus limiting their usability, robustness and generalization. The challenge of data collection becomes particularly pronounced in tasks involving objects with varying or uncertain properties, such as mass or friction. In such cases, substantial and often expensive data acquisition efforts are required to develop a robust policy capable of generalizing effectively across diverse contextual variations [4]. Self-directed exploration in RL, on the other hand, facilitates the development of adaptable strategies that can be generalized across diverse object properties and contexts. These advantages position RL as a compelling approach for learning multisensory contact-rich manipulation [26, 6, 32]. Recent advances in multimodal learning [2, 15] have shown the advantages of reconstructing masked or noised inputs to learn expressive cross-modal representations for downstream tasks. Such masking-based self-supervision approaches, whether at the sensory or embedding level, have also been shown to improve the network’s robustness [45, 31, 41].
@@
To summarize, the key contributions of our work are twofold: (i) we develop an effective pretraining strategy based on masked autoencoding to form a rich multisensory representation, and (ii) we introduce a novel architecture that allows task-specific feature extraction to achieve efficient robot Reinforcement Learning.
Our findings pave the way for adaptive and resilient RL agents to handle multiple input modalities to master complex contact-rich manipulation tasks.
@@ II-A Reinforcement Learning for Contact-Rich Manipulation
Tasks where a robotic manipulator has to interact with its environment, either directly or indirectly via a tool, require a good understanding of the interaction forces that shape the task at hand. Solving contact-rich tasks often relies on an accurate estimate of force and dynamics. When such estimates are available, classical control approaches can be adapted to solve the task [49, 48]. To better handle unseen or difficult contact dynamics in unstructured environments, RL presents itself as an ideal candidate that can learn directly via interactions [40], without accurate state estimation [27]. Various works have employed RL to learn contact-rich manipulation policies on a variety of tasks [10, 46, 33].
@@ II-B Multimodal Self-Supervision for Robotics
Lee et al. [26] learns a multimodal representation from vision, force torque, and proprioception via MLP-fusion and multiple self-supervised objectives. The frozen representation leads to a robust representation for RL to solve multiple peg insertion tasks. [6] extended the Vision Transformer [9] with a force torque sensor to solve a variety of contact-rich tasks. They additionally incorporate the self-supervised objectives from [26] to shape a representation using SLAC [24]. [37] pretrains an audio-encoder [39] to combine vision and audio via a transformer decoder for Imitation Learning. The audio signal from the contact microphone provides rich feedback for various manipulation tasks. To also account for different sensor frequencies [42] developed a multi-resolution policy based on pretrained Vision Language Models to improve inference time using proprioception and force.
@@ II-C Deep Sensor Fusion
A robust representation is essential to integrate and process various sensory inputs to ensure stable and efficient learning in multisensory RL. Previous architectures often focused on straightforward latent fusion approaches by concatenating the various representations for downstream tasks [25, 29, 28, 16, 18]. Feng et al. [11] take this a step further by adding a subgoal-aware weighting for learning the stage-wise importance of difference sensors. Alternatively, contrastive learning approaches [8, 35] emphasize feature alignment, rather than fusion, to learn a shared latent representation between multiple modalities. However, distilling task-relevant features while maintaining sensor-specific information can be challenging. In contrast, inspired by the success of masked token prediction [2], recent works have also explored how masked multisensory pretraining can enhance representation learning for contact-rich manipulation [44, 32].
@@ III Preliminaries
We define a multisensorial POMDP as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{O}^{MS},\ \mathcal{A},\ \mathcal{R},\ \mathcal{P},\ \gamma\rangle$ [23], where $\mathcal{S}$ is the true state of the environment. The agent does not have access to $\mathcal{S}$ and observes it through multiple sensor observations, which we specify as Proprioception $\mathcal{O}^{P}\in\mathbb{R}^{14}$ (joint position and velocities), Force Torque (FT) $\mathcal{O}^{FT}\in\mathbb{R}^{4\times 6}$ (FT-sensor at robot wrist) and Vision $O^{V}\in\mathbb{R}^{64\times 64\times 3}$ (external RGB camera). The sensors build up the multisensory Observation-space $\mathcal{O}^{MS}=[\mathcal{O}^{P},\ \mathcal{O}^{FT},\ \mathcal{O}^{V}]$. $\mathcal{A}$ is the action space (cartesian control), $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is the reward function, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\to\mathcal{S}$ is the transition kernel, and $\gamma\in[0,1)$ is the discount factor. Our goal is to learn a policy $\pi:\mathcal{O}^{MS}\to\mathcal{A}$ that maximizes the discounted sum of rewards $\mathbb{E}_{\pi}[\sum_{t}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$.
We aim to extract an expressive, fused representation from all sensors in order to solve contact-rich manipulation tasks with RL. We only provide the current multisensory observation and no history. We use the off-policy algorithm SAC [17] in simulation and RLPD [3] for efficiency in the real world, however we highlight that our approach is compatible with any actor-critic algorithm.
@@ IV Multisensory Dynamic Pretraining
We present MultiSensory Dynamic Pretraining (MSDP), a novel framework for representation learning that builds upon Masked Autoencoders [20], tailored to enhance contact-rich robot reinforcement learning tasks that require perception through multiple sensor modalities. MSDP introduces a modular architecture to seamlessly handle diverse input sensors, and a multisensory masking scheme to promote rich cross-modal representations, i.e., to retain knowledge about the task and the environment dynamics even in the absence or perturbation of one or more modalities.
Furthermore, we shed light on the different ways to map pretrained transformer embeddings to input states for downstream RL tasks, referred to as latent bridging, a problem often overlooked in practice by the community. In this context, we propose employing a simple cross-attention layer to obtain expressive task-specific features from frozen multisensory embeddings for the critic representation, which is also referred to as attention pooling [5]. The actor on the other hand, receives a special pooling of the sensor embeddings to ensure a stable representation. Both mechanisms offer a fixed low-dimensional observation for the downstream RL agent.
@@ IV-A MSDP Architecture
Given the inputs from multiple sensor modalities, such as vision, force torque readings and proprioception, we encode each modality using a separate network referred to as sensor encoders. Particularly for the vision input, we use a 4 layer CNN-stem similar to [43], which allows us to mask at embedding- rather than pixel-level, as masked object positions cannot be retrieved in pixel-space. The CNN-stem introduces redundancy as vision embeddings have overlapping receptive fields and, most importantly, stabilizes training [52]. Furthermore, compared to a patchify-stem, it reduces the burden of the multisensory encoder to extract vision features, while fusing with other modalities. We use a linear projection to encode force torque readings and proprioception to the 128-dim embedding space given their low dimensionalities.
Subsequently, we incorporate position encoding for the embedded sensor features to contextualize the position and modality of each embedding. Rather than using fixed positional/modality encoding [47] we use learnable embeddings to encode both together. Once we obtain embedding representations from all sensors, we randomly mask out a subset (70 %) of sensor tokens and feed the remaining to our 2 layer transformer encoder with 4 attention heads, pre-normalization and a mlp ratio of 2. The attention mechanism of the encoder leads to dynamic multisensory fusion promoted by the representation objective described in Section IV-B. A learnable mask embedding is added to the randomly masked out embedding positions, next to the multisensory embeddings from the encoder, to generate the original number of embeddings. We reapply our positional/modality encoding before feeding all embeddings to the decoder. During decoding, embeddings exchange information, especially the mask embedding, before being fed to separate decoder heads to fulfill the representation learning objective. We use a shared linear projection for all vision embeddings to reconstruct their corresponding patch.
@@ IV-B Representation Learning
Our representation learning objective is based on multimodal masked autoencoding [20, 2, 15]. The objective is to either reconstruct the current $\mathbf{O}^{MS}_{t}$ or next observation $\mathbf{O}^{MS}_{t+1}$ from a random subset of sensor embeddings. Vision has a high number of embeddings and needs to extract information about other, potentially masked, sensors e.g. identifying contact to estimate force. Non-vision sensors on the other hand, are beneficial to reconstruct the vision observation as e.g. proprioception defines the robot position. Vision as a global sensing modality never gets fully masked out, whereas low-dimensional sensors are not available when masked. This representation objective results in cross-sensor prediction thus leading to fusion of all modalities. Furthermore, the decoder is conditioned on action in the prediction objective and needs to extract dynamic action-related features important for downstream RL. We denote our method MSDP-P (Prediction) wherein we reconstruct the next observation $\mathbf{O}^{MS}_{t+1}-\Phi(\mathbf{O}^{MS}_{t},\mathbf{A}_{t})||^{2}$ and MSDP-R (Reconstruction) where we reconstruct the current observation $\mathbf{O}^{MS}_{t}-\Phi(\mathbf{O}^{MS}_{t})||^{2}$. $\Phi(.)$ denotes the network prediction conditioned on the current observation $\mathbf{O}^{MS}_{t}$ and action $\mathbf{A}_{t}$. $||\cdot||^{2}$ denotes the Mean Squared Error. Sensor embeddings are only masked during pretraining.
@@ IV-C Policy Learning
From the pretrained encoder, we receive expressive multisensory embeddings given the available sensors. Extracting a compact representation from those embeddings for downstream task-solving is a crucial and often under-discussed design choice. We name the mapping between embeddings and a compact representation latent bridging, which has a considerable impact on performance. To address this, we depart from works that naively extract the "CLS"-token from the high-dimensional embeddings, analogous to the Vision Transformer [41, 9, 51, 36] and, instead, propose an asymmetric latent bridging strategy between actor and critic to obtain a compact representation. The critic uses a single cross-attention layer with a learnable query and the multisensory embeddings from the MSDP encoder as keys and values (see Figure 2). It offers dynamic task-specific feature extraction (e.g. object positions, robot state, contact) over the task-solving process. Fine-grained understanding of the environment leads to faster convergence compared to a global representation. The policy, on the other hand, does not profit from a cross-attention layer as it may destabilize training. Instead, we mean pool all embeddings originating from the vision sensor before pooling all embeddings, to account for the uneven number of sensor embeddings. Pooling sensor tokens results in a stable and parameter-free latent bridging similar to [44, 43]. This assymmetric representation between actor and critic, follows [14], where the actor benefits from a stable representation over task stages and the critic from a dynamic-specific representation. The cross-attention layer is trained by the critic, resulting in task-specific feature extraction. We train with the off-policy algorithm SAC in simulation and RLPD in the real world, where we incorporate offline data for pretraining in our replay buffer. For pretraining and policy training we use a batch size of 64 and a learning rate of $3\times 10^{-4}$. Our approach is working with any actor-critic RL algorithm.